Describe the preparation of the following solution (show all calculations): 5.0 L of 6.000 M H3PO4 from a concentrated solution that is 85.1%(m/m) and has a specific gravity of 1.69
1
Expert's answer
2020-06-15T14:43:01-0400

In order to prepare "V"= 5 L of "c" = 6.000 M H3PO4 solution, one needs the following quantity "n" of H3PO4:

"n(H_3PO_4) = cV=5\u00b76.000 = 30" mol.

The mass of this quantity of H3PO4 can be calculated using the molar mass "M" of H3PO4 (98.00 g/mol):

"m = nM = 30\u00b798.00 =2940" g.

Thus, the mass of the concentrated solution needed contains 2940 g of H3PO4:

"m_s = \frac{2940}{0.851} = 3454.8" g.

The specific gravity of the concentrated solution is 1.69 g/mL. Therefore, 3454.8 g of the concentrated solution corresponds to the following volume:

"V_s = \frac{m}{d} = \frac{3454.8}{1.69}=2044" mL.

Then, to 2044 mL of the concentrated solution one needs to add the following volume of water:

"V_w = 5000-2044 = 2956" mL.

Finally, in order to prepare 5 L of 6.000 M of H3PO4 from a concentrated solution

that is 85.1%(m/m) and has a specific gravity of 1.69 g/mL, one needs to take 2044 mL of the concentrated solution and add 2956 mL of water.